OpenStudy (anonymous):

decide whether the congruent statement is true or false 4=9 (mod 2)

OpenStudy (zzr0ck3r):

does 2 divide 4-9?

OpenStudy (zzr0ck3r):

@racheltat ?

OpenStudy (anonymous):

no! @zzr0ck3r so does that mean its false? do you just subtract the two numbers?

OpenStudy (zzr0ck3r):

correct \(a\equiv b\text{ mod}(c)\implies c|(a-b)\)

OpenStudy (zzr0ck3r):

\(a\equiv b\text{ mod}(c)\iff c|(a-b)\)

OpenStudy (zzr0ck3r):

it goes both ways....

OpenStudy (anonymous):

wow thanks! so 1106=551 (mod 5 ) would be true because 1106-551 equals 555 and that goes in 5?

OpenStudy (zzr0ck3r):

5 goes into 555

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

awesome!! sorry to bug you but it seems like you understand mod arithmetic well and i dont lol... example: (7x10) (mod 6) how would i solve a problem like that?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

Im think the answer would be 6?

OpenStudy (zzr0ck3r):

it just wants the remainder of 70 when divided by 6

OpenStudy (zzr0ck3r):

so 4

OpenStudy (anonymous):

oh gotcha so 11! wow thank you so much

OpenStudy (zzr0ck3r):

11?

OpenStudy (anonymous):

oops never mind i was looking at different problem lol.. i have many equations lol.. thank you for your help!!

OpenStudy (zzr0ck3r):

70 = 11*6+4, the answer is the 4

OpenStudy (zzr0ck3r):

word, np

OpenStudy (anonymous):

wait one more thing lol! (43-23) (mod 5) would be 4 right? 43-23 = 20 20/5= 4?

OpenStudy (anonymous):

@zzr0ck3r

ganeshie8 (ganeshie8):

try again :) you need to stare at "remainder", not "quotient"

OpenStudy (zzr0ck3r):

0

OpenStudy (zzr0ck3r):

43-23 = 20, and for sure 5 divides 20, so it has a remainder of 0

OpenStudy (zzr0ck3r):

these questions are simply asking for the remainder of the number when it is divided by the thing in the mod()

OpenStudy (zzr0ck3r):

@ganeshie8 what do we call the mod(*) thing? not an operator is it?

OpenStudy (anonymous):

oh!! ok i understand! I'm trying to figure out whats left over! that makes so much sense. thank you!! @zzr0ck3r @ganeshie8